# Human Cognition Inspired RAG with Knowledge Graph for Complex Problem Solving

### v1

@@ Appendix B Baselines
In this subsection, we introduce the baselines used in our experiments, including LLaMA2-13B, LLaMA3-8B, Qwen2.5-7B, Chain-of-Thought (CoT) prompting, Think-on-graph (ToG), MindMap and Graph-CoT.
@@
LLaMA2-13B Touvron et al. (2023) is a member of the LLM series developed by Meta and is the second generation version of the LLaMA (Large Language Model Meta AI) model. LLaMA2-13B contains 13 billion parameters and is a medium-sized large language model.
@@
LLaMA3-8B AI@Meta (2024) is an efficient and lightweight LLM with 8 billion parameters, designed to provide high-performance natural language processing capabilities while reducing computing resource requirements.
@@
Qwen2.5-7B Yang et al. (2024) is an efficient and lightweight LLM launched by Alibaba with 7 billion parameters. It aims at making it more efficient, specialized, and optimized for particular applications.
@@
Chain-of-Thought Wei et al. (2022) is a method designed to enhance a model’s reasoning ability, particularly for complex tasks that require multiple reasoning steps. It works by prompting the model to generate intermediate reasoning steps rather than just providing the final answer, thereby improving the model’s ability to reason through complex problems.
@@
Think-on-graph Sun et al. (2023) proposed a new LLM-KG integration paradigm, “LLM $\otimes$ KG”, which treats the LLM as an agent to interactively explore relevant entities and relationships on the KG and perform reasoning based on the retrieved knowledge. It further implements this paradigm by introducing a novel approach called “Graph-based Thinking”, in which the LLM agent iteratively performs beam search on the KG, discovers the most promising reasoning paths, and returns the most likely reasoning result.
@@
MindMap Wen et al. (2023) propose a novel prompting pipeline, that leverages KGs to enhance LLMs’ inference and transparency. It enables LLMs to comprehend KG inputs and infer with a combination of implicit and external knowledge. Moreover, it elicits the mind map of LLMs, which reveals their reasoning pathways based on the ontology of knowledge.
@@
Graph-CoT Jin et al. (2024b) propose a simple and effective framework called Graph Chain-of-thought to augment LLMs with graphs by encouraging LLMs to reason on the graph iteratively. Moreover, it manually construct a Graph Reasoning Benchmark dataset called GRBENCH, containing 1,740 questions that can be answered with the knowledge from 10 domain graphs.
Figure 5: Case of CogGRAG.
@@ Appendix C Case Studies
Appendix D Prompts in CogGRAG
Table 6: The prompt template for decomposition.
Decomposition prompt
Prompt head: “Your task is to decompose the given question Q into sub-questions.
You should based on the specific logic of the question to determine the number of sub-questions
and output them sequentially. ”
Instruction: “Please only output the decomposed sub-questions as a string in list format, where each
element represents the text of a sub-question, in the form of ’[\"subq1\", \"subq2\", \"subq3\"]’.
For each sub-questions, if you consider the sub-question to be sufficiently simple and no
further decomposition is needed, then output \"End.\", otherwise, output \"Continue.\".
Please strictly follow the format of the example below when answering the question.
Here are some examples: ”
“ Input: “What year did Guns N Roses perform a promo for a movie starring Arnold Schwarzenegger
as a former New York Police detective?”
Output: [ { “Sub-question”: “What movie starring Arnold Schwarzenegger as a former New York Police
detective is being referred to?”,
“State”: “Continue.” },
{ “Sub-question”: “In what year did Guns N Roses perform a promo for the movie mentioned
in sub-question #1?”,
“State”: “End.” } ]
Input: “What is the name of the fight song of the university whose main campus is in Lawrence, Kansas
and whose branch campuses are in the Kansas City metropolitan area?”
Output: [ { “Sub-question”: “Which university has its main campus in Lawrence, Kansas and branch
campuses in the Kansas City metropolitan area?”,
“State”: “End.”
{ “Sub-question”: “What is the name of the fight song of the university identified in sub-question #1?”,
“State”: “End.” } ]
Input: “Are the Laleli Mosque and Esma Sultan Mansion located in the same neighborhood?”
Output: [ { “Sub-question”: “Where is the Laleli Mosque located?”,
“State”: “End.”
{ “Sub-question”: “Where is the Esma Sultan Mansion located?”,
“State”: “End.” },
{ “Sub-question”: “Are the locations of the Laleli Mosque and the Esma Sultan Mansion in the same
neighborhood?”,
“State”: “End.” } ] ”
Input: Question Q
Output:
Table 7: The prompt template for extraction on local level.
Extraction prompt
Prompt head: “Your task is to extract the entities (such as people, places, organizations, etc.)
and relations (representing behaviors or properties between entities, such as verbs,
attributes, or categories, etc.) involved in the input questions. These entities and relations can
help answer the input questions.”,
Instruction: “Please extract entities and relations in one of the following forms: entity, tuples,
or triples from the given input List. Entity means that only an entity, i.e. <entity>. Tuples means that
an entity and a relation, i.e. <entity-relation>. Triples means that complete triples, i.e. <entity-relation-entity>.
Please strictly follow the format of the example below when answering the question. ”,
Input: [The mind map $M$]
Output:
Table 8: The prompt template for extraction on global level.
Extraction prompt
Prompt head: “Your task is to extract the subgraphs involved in a set of input questions.”,
Instruction: “Please extract and organize information from a set of input questions into structured
subgraphs. Each subgraph represents a group of triples (subject, relation, object) that share common
entities and capture the logical relationships between the questions. Here are some examples:”
“ Input: [“What is the capital of France?”, “Who is the president of France?”,
“What is the population of Paris?”]
Output: [(“France”, “capital”, “Paris”), (“France, “president”, “Current President”),
(“Paris”, “population”, “Population Number”) ] ”
Input: [The mind map $M$]
Output:
Table 9: The prompt template for reasoning.
Reasoning prompt
Prompt head: “Your task is to answer the questions with the provided completed reasoning and input
knowledge.”,
The completed reasoning: [The reasoning process $\hat{M}$]
The knowledge graph: [Knowledge $\mathcal{T}$]
Input: [Subquestion $q^{t}_{j}$]
Output:
Table 10: The prompt template for self-verification.
Self-verification prompt
Prompt head: “You are a logical verification assistant. Your task is to check whether the answer to a
given question is logically consistent with the provided completed reasoning and input knowledge.
If the answer is consistent, respond with “right”. If the answer is inconsistent, respond with “wrong”.”
The completed reasoning: [The reasoning process $\hat{M}$]
The knowledge graph: [Knowledge $\mathcal{T}$]
Input: [Subquestion $q^{t}_{j}$]
Answer: [Answer $a^{t}_{j}$]
Output:
Table 11: The prompt template for re-thinking.
Re-thinking prompt
Prompt head:“You are a reasoning and knowledge integration assistant. Your task is to re-think
a question that was previously answered incorrectly by the self-verification model. Use the provided
completed reasoning and input knowledge to generate a new answer.”
Instruction: “Please note, if the knowledge is insufficient to answer the question, respond
with “Insufficient information, I don’t know”.The response must be included in square brackets [xxx].”
The completed reasoning: [The reasoning process $\hat{M}$]
The knowledge graph: [Knowledge $\mathcal{T}$]
Input: [Subquestion $q^{t}_{j}$]
Output:

### v2

@@ Appendix B Baselines
In this subsection, we introduce the baselines used in our experiments, including LLaMA2-13B, LLaMA3-8B, Qwen2.5-7B, Qwen2.5-32B, Chain-of-Thought (CoT) prompting, Think-on-graph (ToG), MindMap, Graph-CoT, Reasoning on Graphs (RoG), Generate-on-Graph (GoG).
@@
LLaMA2-13B (llama2) is a member of the LLM series developed by Meta and is the second generation version of the LLaMA (Large Language Model Meta AI) model. LLaMA2-13B contains 13 billion parameters and is a medium-sized large language model.
@@
LLaMA3-8B (llama3) is an efficient and lightweight LLM with 8 billion parameters, designed to provide high-performance natural language processing capabilities while reducing computing resource requirements.
@@
Qwen2.5-7B (qwen2) is an efficient and lightweight LLM launched by Alibaba with 7 billion parameters. It aims at making it more efficient, specialized, and optimized for particular applications.
Qwen2.5-32B (hui2024qwen2) is an efficient LLM launched by Alibaba with 32 billion parameters. It is a more comprehensive foundation for real-world applications such as Code Agents. Not only enhancing coding capabilities but also maintaining its strengths in mathematics and general competencies.
@@
Chain-of-Thought (prompt_cot) is a method designed to enhance a model’s reasoning ability, particularly for complex tasks that require multiple reasoning steps. It works by prompting the model to generate intermediate reasoning steps rather than just providing the final answer, thereby improving the model’s ability to reason through complex problems.
@@
Think-on-graph (Think_on_graph) proposed a new LLM-KG integration paradigm, “LLM $\otimes$ KG”, which treats the LLM as an agent to interactively explore relevant entities and relationships on the KG and perform reasoning based on the retrieved knowledge. It further implements this paradigm by introducing a novel approach called “Graph-based Thinking”, in which the LLM agent iteratively performs beam search on the KG, discovers the most promising reasoning paths, and returns the most likely reasoning result.
@@
MindMap (mindmap) propose a novel prompting pipeline, that leverages KGs to enhance LLMs’ inference and transparency. It enables LLMs to comprehend KG inputs and infer with a combination of implicit and external knowledge. Moreover, it elicits the mind map of LLMs, which reveals their reasoning pathways based on the ontology of knowledge.
@@
Graph-CoT (graph_cot) propose a simple and effective framework called Graph Chain-of-thought to augment LLMs with graphs by encouraging LLMs to reason on the graph iteratively. Moreover, it manually construct a Graph Reasoning Benchmark dataset called GRBENCH, containing 1,740 questions that can be answered with the knowledge from 10 domain graphs.
Reasoning on Graphs (luo2024rog) propose a novel method called reasoning on graphs (RoG) that synergizes LLMs with KGs to enable faithful and interpretable reasoning. Specifically, they present a planning retrieval-reasoning framework, where RoG first generates relation paths grounded by KGs as faithful plans. These plans are then used to retrieve valid reasoning paths from the KGs for LLMs to conduct faithful reasoning.
Generate-on-Graph (xu2024gog) propose leveraging LLMs for QA under Incomplete Knowledge Graph (IKGQA), where the provided KG lacks some of the factual triples for each question, and construct corresponding datasets. To handle IKGQA, they propose a training-free method called Generate-on-Graph (GoG), which can generate new factual triples while exploring KGs.
@@ Appendix C Case Studies
Appendix D Prompts in CogGRAG
$\texttt{prompt}_{dec}$
Prompt head: “Your task is to decompose the given question Q into sub-questions. You should based on the specific logic of the question to determine the number of sub-questions and output them sequentially.”
Instruction: “Please only output the decomposed sub-questions as a string in list format, where each element represents the text of a sub-question, in the form of {[“subq1”, “subq2”, “subq3”]}. For each sub-question, if you consider the sub-question to be sufficiently simple and no further decomposition is needed, then output “End.”, otherwise, output “Continue.” Please strictly follow the format of the example below when answering the question.
Here are some examples:”
“Input: “What year did Guns N Roses perform a promo for a movie starring Arnold Schwarzenegger as a former New York Police detective?”
Output: [ { “Sub-question”: “What movie starring Arnold Schwarzenegger as a former New York Police detective is being referred to?”, “State”: “Continue.” }, { “Sub-question”: “In what year did Guns N Roses perform a promo for the movie mentioned in sub-question #1?”, “State”: “End.” } ]”
“Input: “What is the name of the fight song of the university whose main campus is in Lawrence, Kansas and whose branch campuses are in the Kansas City metropolitan area?”
Output: [ { “Sub-question”: “Which university has its main campus in Lawrence, Kansas and branch campuses in the Kansas City metropolitan area?”, “State”: “End.” }, { “Sub-question”: “What is the name of the fight song of the university identified in sub-question #1?”, “State”: “End.” } ]”
“Input: “Are the Laleli Mosque and Esma Sultan Mansion located in the same neighborhood?”
Output: [ { “Sub-question”: “Where is the Laleli Mosque located?”, “State”: “End.” }, { “Sub-question”: “Where is the Esma Sultan Mansion located?”, “State”: “End.” }, { “Sub-question”: “Are the locations of the Laleli Mosque and the Esma Sultan Mansion in the same neighborhood?”, “State”: “End.” } ]”
Input: Question $Q$
Output:
Table 7: The prompt template for decomposition.
$\texttt{prompt}_{ext}$
Prompt head: “Your task is to extract the entities (such as people, places, organizations, etc.) and relations (representing behaviors or properties between entities, such as verbs, attributes, or categories, etc.) involved in the input questions. These entities and relations can help answer the input questions.”
Instruction: “Please extract entities and relations in one of the following forms: entity, tuples, or triples from the given input List. Entity means that only an entity, i.e. <entity>. Tuples means that an entity and a relation, i.e. <entity-relation>. Triples means that complete triples, i.e. <entity-relation-entity>. Please strictly follow the format of the example below when answering the question.”
Input: [The mind map $\mathcal{M}$]
Output:
Table 8: The prompt template for extraction on local level.
$\texttt{prompt}_{ext}$
Prompt head: “Your task is to extract the subgraphs involved in a set of input questions.”
Instruction: “Please extract and organize information from a set of input questions into structured subgraphs. Each subgraph represents a group of triples (subject, relation, object) that share common entities and capture the logical relationships between the questions. Here are some examples:”
“Input: [“What is the capital of France?”, “Who is the president of France?”, “What is the population of Paris?”]
Output: [(“France”, “capital”, “Paris”), (“France”, “president”, “Current President”), (“Paris”, “population”, “Population Number”)]”
Input: [The mind map $\mathcal{M}$]
Output:
Table 9: The prompt template for extraction on global level.
$\texttt{prompt}_{res}$
Prompt head: “Your task is to answer the questions with the provided completed reasoning and input knowledge.”
The completed reasoning: [The set of verified answers $\hat{\mathcal{M}}$]
The knowledge graph: [The final retrieved triple set $\mathcal{T}$]
Input: [Subquestion $q^{t}$]
Output:
Table 10: The prompt template for reasoning.
$\texttt{prompt}_{ver}$
Prompt head: “You are a logical verification assistant. Your task is to check whether the answer to a given question is logically consistent with the provided completed reasoning and input knowledge. If the answer is consistent, respond with “right”. If the answer is inconsistent, respond with “wrong”.”
The completed reasoning: [The set of verified answers $\hat{\mathcal{M}}$]
The knowledge graph: [The final retrieved triple set $\mathcal{T}$]
Answer: [The candidate answer $a^{t}$]
Input: [Subquestion $q^{t}$]
Output:
Table 11: The prompt template for self-verification.
$\texttt{prompt}_{rethink}$
Prompt head: “You are a reasoning and knowledge integration assistant. Your task is to re-think a question that was previously answered incorrectly by the self-verification model. Use the provided completed reasoning and input knowledge to generate a new answer.”
Instruction: “Please note, if the knowledge is insufficient to answer the question, respond with “Insufficient information, I don’t know”. The response must be included in square brackets [xxx].”
The completed reasoning: [The set of verified answers $\hat{\mathcal{M}}$]
The knowledge graph: [The final retrieved triple set $\mathcal{T}$]
Input: [Subquestion $q^{t}$]
Output:
Table 12: The prompt template for re-thinking.